Use the Henderson-Hasselbalch equation to determine the ratio of acid to base in a formic acid-formate...

  1. Use the Henderson-Hasselbalch equation to determine the ratio of acid to base in a formic acid-formate buffer with a pH of 3.00

  2. Assume that the human blood buffer includes 0.00080 M carbonic acid and 0.0080 M hydrogen carbonate. What is the pH of 7.00 liter of blood under these conditions?

  3. How many moles of strong acid must be added to the blood in question 2 to bring the hydrogen carbonate/ carbonic acid ratio to the hazardous level of 5:1 AND what would the pH be?

  4. How many moles of hydrogen carbonate need to be added to the blood in question 3 to raise the pH to 7.45?
Answer #1

1)- Ratio of Acid to base is 5.62.A

2)- pH of blood at the given conditions is 7.35.

3)- Moles of strong acid added = 3.115*10-7

4)- Amount of hydrogen carbonate added = 0.014 molAccording to Henderson - Hasselbalch equation. pH a pka + los CA71 [HA] , Where , Ka - Acid dissociation constant [HA] - conc0.83 Here the human blood is a carbonic acid and its conjugate 50, Handerson - Hasselbalch qqn buffer of base is valid CAL 5)In ay Here, [A]. 5. [HAT . 2) PH 6.35 + log10 (5) 5) PH = 6.35 +0.7 = 7,05 a) JPH 3.7:05 ) concentration of H= 10 17:05 GritiDay 7 Here, PHE PKO + Cogno con pH = pka = 7.45 6.35 6) =) [A] - 12.56 X 0.0008 = 0.01 = concentration of Base Jiras = orod g